OpenStudy (anonymous):

Find the percent of discount; regular price:$200 discount;$30

OpenStudy (anonymous):

60

OpenStudy (anonymous):

Anytime

OpenStudy (anonymous):

can you please show how u got this answer?

OpenStudy (anonymous):

Is $30 the discounted price or it was $30 off?

OpenStudy (anonymous):

Multiply and than divide.

OpenStudy (anonymous):

Use a calculator man

OpenStudy (anonymous):

okay thanks, and it says its the sale price

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

If original price is $200 and sale price is $30, discount percent = (200-30)/200 = 170/200 = 85%

OpenStudy (anonymous):

GT has the right answer

OpenStudy (anonymous):

X is percentage discount 200- 200*x= 30 200(1 -x) = 30 1-x= 30/200 1-30/200 = x 1-3/20= x 20/20 - 3/20 =x 17/20=x .85 = x